Suppose that an algorithm has run-time proportional to 2n , where n is the input size. The algorithm takes 1 millisecond to process an array of size 10. How many milliseconds would you expect the algorithm take to process an array of size 20 ?
Answer #1

1)

algorithm has runtime of 2

\frac{t1}{t2}=\frac{n1}{n2}

\frac{t1}{t2}=\frac{2^{20}}{2^{10}}=2^{10}=1024

so,\;t1=1024*t2

given\;t2=1millisecond

t1=1024\;milliseconds=1.024\;seconds

\color{Red}Answer:=1.024\;seconds

Note:\;\;if\;runtime\;is\;2n\;instead\;of\;2^n\;then\;answer\;is\;2\;milliseconds.